### v1

# C3VDv2 - Colonoscopy 3D video dataset with enhanced realism

Abstract
Computer vision techniques have the potential to improve the diagnostic performance of colonoscopy, but the lack of 3D colonoscopy datasets for training and validation hinders their development. This paper introduces C3VDv2, the second version (v2) of the high-definition Colonoscopy 3D Video Dataset, featuring enhanced realism designed to facilitate the quantitative evaluation of 3D colon reconstruction algorithms. 192 video sequences were captured by imaging 60 unique, high-fidelity silicone colon phantom segments. Ground truth depth, surface normals, optical flow, occlusion, six-degree-of-freedom pose, coverage maps, and 3D models are provided for 169 colonoscopy videos. Eight simulated screening colonoscopy videos acquired by a gastroenterologist are provided with ground truth poses. The dataset includes 15 videos featuring colon deformations for qualitative assessment. C3VDv2 emulates diverse and challenging scenarios for 3D reconstruction algorithms, including fecal debris, mucous pools, blood, debris obscuring the colonoscope lens, en-face views, and fast camera motion. The enhanced realism of C3VDv2 will allow for more robust and representative development and evaluation of 3D reconstruction algorithms.
Background & Summary
Colorectal cancer (CRC) is a significant public health concern, ranking as the second leading cause of cancer-related deaths in the United States [1]. Colonoscopy is the gold standard procedure for identifying and resecting precancerous lesions in the gastrointestinal tract [2]. The timely detection and removal of precancerous lesions through routine colonoscopy screening is estimated to reduce the CRC mortality rate by 53%[3]. The protective value of screening colonoscopy depends on the operator’s ability to thoroughly inspect the full colon surface and identify all clinically relevant lesions. A meta-analysis of over 15,000 tandem colonoscopies revealed that approximately 26% of adenomas go undetected [4]. In a retrospective analysis, McGill et al. [5] estimated that 19% of the colon surface area typically remains unobserved during a colonoscopy, potentially leading to missed lesions.
3D computer vision algorithms, including Simultaneous Localization and Mapping (SLAM), Neural Radiance Fields (NeRFs), and 3D Gaussian Splatting, have been used to reconstruct the colon surface from colonoscopy videos [6, 7, 8, 9]. Due to the enclosed nature of the colon, unobserved regions appear as ‘gaps’ in 3D reconstructions, which could highlight areas missed by clinicians [10]. Real-time systems employing these techniques could assist navigation during colonoscopies and increase mucosal coverage. Furthermore, advanced quality metrics, like fractional coverage area, can better assess screening effectiveness. SLAM algorithms include submodules such as camera tracking [11, 12] and depth estimation [13, 14]. These submodules also enable downstream applications such as anatomical landmark recognition [15], using depth to improve lesion classification [16], polyp size estimation [17], and augmented reality for training novice gastroenterologists [18].
3D reconstruction of the colon from colonoscopy video data presents several important challenges for computer vision algorithms. These include the absence of distinct and repetitive features due to the textureless appearance of the colon surface, specular reflections, soft tissue deformation from peristalsis or distension, and occlusions caused by haustral folds [19]. Additional difficulties arise from the presence of fecal debris, rapid scope motion, limited viewpoints within the constrained tubular anatomy, wide range of working distances present in many scenes, and non-uniform illumination from the co-located light source and camera. Despite their clinical relevance, such challenging factors are rarely captured together in existing datasets that provide quantitative 3D ground truth. Thus, there is a need for datasets providing pixel-wise 3D ground truth for realistic and challenging colonoscopy scenarios.
There are limited options for obtaining in vivo colon 3D ground truth information. Current monocular colonoscopy datasets derive 3D information from Computed Tomography (CT) scans of the human colon [20, 21, 22, 23], and optical scanning of ex vivo animal tissue [24]. Each approach offers a trade-off between realism and ground truth accuracy. Although datasets imaging ex vivo [24] or in vivo tissue [25] provide realistic videos, they lack pixel-wise 3D ground truth. Conversely, simulated datasets generated with rendering engines offer accurate pixel-level ground truth but struggle to replicate real-world camera optics, illumination models, and light-tissue interactions [23, 26, 27, 20, 24, 21]. Moreover, most synthetic datasets fail to incorporate realistic conditions such as fecal debris, blood, colon deformations, and the presence of endoscopy instruments. A recent synthetic dataset has made progress in this area by providing realistic vasculature, polyps, ulcers, and blood simulations, but it primarily focuses on lesion segmentation and classification, and thus lacks 3D ground truth [28].
To address these limitations, Bobrow et al. [29] introduced the Colonoscopy 3D Video Dataset (C3VD), which utilizes high-fidelity silicone phantoms with known 3D shapes, imaged with a high-definition clinical colonoscope. C3VD’s primary advantage over real colonoscope datasets, such as EndoSLAM [24] and EndoMapper [25], is providing per-frame pixel-level registered ground truth data. By imaging phantoms with a clinical colonoscope, C3VD inherently captures non-global illumination, real fisheye intrinsics, and non-linear post-processing effects, thereby addressing key limitations of simulated datasets [23, 20, 24, 21]. However, the original C3VD videos only spanned idealized scenarios with clean colons and lenses, slow and smooth camera motion, a static colon, and no tools and accessories present in the field of view (FoV).
C3VDv2 retains the advantages of C3VD, including clinical HD colonoscope videos with registered ground pixel-wise depth maps, normals, optical flow, occlusion maps, diffuse images, and camera poses. Additionally, it introduces several enhancements over the original C3VD dataset to improve realism and comprehensiveness, and increases the dataset size. More specifically, the C3VDv2 dataset incorporates the following improvements:
  1. 1.
    Larger dataset: The number of videos has increased eightfold (n=192). The variety of colon geometries has expanded twofold, with eight full colon phantoms.
  2. 2.
    Realistic artifacts: The realism of colon appearance has been significantly improved by including artifacts such as fecal debris, mucus pools, blood, foam, and debris and water on the lens. Furthermore, C3VDv2 incorporates instruments like water jets, lens cleaning, and suction.
  3. 3.
    Challenging scenarios: The dataset also includes scenarios that are particularly challenging for 3D reconstruction and tracking algorithms. These include fast and abrupt camera motion, en-face to down-the-barrel pose transitions, close-up en-face views of textureless surfaces, the endoscope getting covered in debris, and lens cleaning. The dataset also includes trajectories such as straight-line in-and-out motions, loops where the first and last points are the same, and paths where the first half mirrors the second half with lens cleaning in the middle.
  4. 4.
    Colon deformation: C3VDv2 contains videos of colon deformation. Due to methodological limitations, these videos lack pixel-wise ground truth maps, but they do include camera poses and the ground truth of the undeformed 3D models, making it a valuable resource for qualitative assessment or quantitative assessment after a surface transformation (“unstretching”).
  5. 5.
    Paired clean & debris-filled colon frames: Lastly, for every debris-filled colon video, there is a corresponding clean colon video acquired by imaging a colon phantom with no debris using an identical camera trajectory as the debris-filled video. These paired videos enable the quantification of performance degradation due to debris.
C3VDv2 is a valuable resource for training and validating endoscopy 3D reconstruction algorithms. The dataset is publicly available and includes all 3D models, camera pose trajectories, coverage mesh, and occlusion maps, essential for validating SLAM algorithms to identify missing regions. Pixel-level depth estimation, surface normals, and optical flow data can be used to validate both deep learning algorithms and components of SLAM methods. Relative to the first version of C3VD, C3VDv2 dramatically expands the diversity of realistic scenes while increasing the dataset size, enabling comprehensive quantitative evaluation of computer vision algorithms across a wide range of realistic and challenging colonoscopy scenarios.
Methods
The C3VDv2 dataset includes colonoscopy videos with pixel-level ground truth depth maps. To achieve this, we first established a protocol for fabricating high-fidelity phantoms from a 3D surface model of a human colon. Then, we record video sequences of a colonoscope advanced through colon phantoms with a robotic arm while recording the pose log. Finally, we used a multimodal 2D/3D video registration technique to align the acquired video and trajectory data with a ground truth 3D surface model of the colon. A more detailed description of the registration method is described in C3VD [29].
Silicone colon phantom fabrication
To develop realistic colon phantoms for colonoscopy research, two complete 3D colon models were sculpted using MeshMixer software based on colonoscopy videos and the C3VD model. Each model was reviewed by an experienced gastroenterologist for anatomical accuracy. Colon 1 (c1) was divided into eight distinct regions, while Colon 2 (c2) comprised seven regions, aligning with anatomical sections - cecum, ascending colon, transverse colon, descending colon, sigmoid colon, and rectum; further subdivisions such as transverse 1 and transverse 2 were created for longer segments to fit the 3D printer’s build volume (14.5 × 14.5 × 19.3 cm). Lesions representing various categories of the Paris Classification [30] were sculpted onto the models and strategically placed such that the haustral folds would occlude some polyps. Comprehensive details regarding the lengths of model segments, types of lesions, and their major axis diameters are presented in Tables 2 and 3, and an illustration is presented in Figure 3.
Three-part molds were created for each colon segment. The molds comprised two outer parts forming the phantom’s external shell and one insert part forming the colon lumen. All insert parts were 3D printed on a high-resolution Formlabs Form 3 printer using the adaptive layer thickness setting, which used a layer height between 25 and 200 micrometers to depict the colon lumen details accurately. The outer shells, which form the external surface not seen by the colonoscope, were printed using an Ultimaker S3 printer with a 0.2-millimeter resolution.
Silicone (Dragon Skin, Smooth-On, Inc.) was used to create casts from each mold, with color and texture variations introduced using silicone pigments (Silc Pig, Smooth-On, Inc.). For each phantom, the initial layer of silicone was manually layered onto the insert. 5 to 12 layers were applied to allow for detailed hand-painting of features such as vasculature patterns on intermediate layers. Silicone was later poured into the bulk of the mold through inlets in the outer molds, filling any gaps between the insert and outer shells. This entire process is illustrated in Figure 2.
Using the molds for each model segment, we cast four complete phantoms, each varying in texture and color. This effort resulted in a total of eight full-length colon phantoms as shown in Figure 4. The diversity in tissue colors and textures among these phantoms was explicitly designed to enable domain randomization for robust training and evaluation of computer vision algorithms.
Data Acquisition
Imaging Setup
The experimental setup employed for data acquisition is illustrated in Figure 2. We used an Olympus CF-HQ190L video colonoscope, CV-190 video processor, and CLV-190 light source to image the phantom models. Uncompressed video frames were captured from the clinical video processor using an Orion HD (Matrox Imaging) graphics adapter. Video frames were parsed and stored in raw uncompressed RGB format. Silicone models were placed inside the outer shell molds to prevent phantom motion and deformation during video recording. All videos were recorded using white light illumination. The colonoscope’s distal tip was rigidly and securely mounted to the end effector of a Universal Robotics UR-3e robotic arm. 6-12 waypoints were manually programmed for each video segment to simulate colonoscopy trajectories. The robotic arm traversed the colon by interpolating between these poses, with a repeatability of ±$\pm 0.03$ mm. A pose log was recorded from the arm at a sampling rate of 500 Hz.
Enhanced Realism
Colonoscopy artifacts, such as specular reflections, mucous pools, fecal debris, blood, and foam, pose significant challenges for 3D reconstruction algorithms. We introduce these artifacts to increase the realism of the colonoscopy videos. Real colonoscopy videos served as a reference for creating the debris-filled colon environment. To emulate the specular reflections of the mucosa, the phantom was coated with a silicone lubricant (015594011516, BioFilm, Inc.). A mixture of peanut butter (051500255377, Jif), applesauce (085239045145, Good & Gather), and yellow pigment powder (B0B82NLDDR, Rock N Soil) mimicked fecal debris. Different grades of bowel preparation were created by varying the quantities of debris. Simulated blood was created by combining glycerine (754207111907, Raw Plus Rare), red dye (B089HYVC3R, Craft County), and applesauce. Mucous pools were simulated using a mixture of glycerine, yellow dye, and chicken soup (051000134592, Campbell). Soap foam, colored with yellow pigment powder or red dye, was also added to the debris. The colonoscope’s auxiliary water jet system was used to flush debris from the phantom’s walls. Water-on-lens and lens cleaning scenarios were captured using the colonoscope’s lens cleaning system. In simulated screening videos, the system’s suction feature was used to drain mucous pools and debris. Representative examples are shown in Figure 1 and 5. These debris formulations were reviewed by an experienced gastroenterologist for qualitative agreement with clinical scenes.
Pixel-level GT Video Collection Protocol
For each phantom, a maximum of four videos were recorded. The first and second videos (v1 & v2) captured the clean colon, exhibiting only specular reflection and occasionally lens cleaning. These videos had different imaging settings, such as illumination intensity, and trajectories. Debris was then added to the phantom for the third video (v3), with the camera trajectory from v2 being replicated. In a few cases, a colon deformation video was also captured (v4). The phantom was removed from the molds and deformed using a blood pressure cuff (31191721394, Walgreens) inflated using an air pump (Pacum, Masterspace). The deformation video (v4) features either a stationary camera or linear camera motion.
Simulated Screening Colonoscopy
In addition to the short videos featuring phantom segments, we created simulated screening colonoscopy videos of complete colon phantoms. Individual phantom segments were bonded together using a silicone adhesive (Sil-Poxy™, Smooth-On, Inc.) and selectively sutured for structural reinforcement. As discussed previously, materials such as glycerin and yellow dye were incorporated to simulate debris. The full colon phantom model was mounted on a cutout foam scaffold to reduce global movement and colon deformation during imaging.
For camera tracking in this experiment, a six-degree-of-freedom electromagnetic (EM) sensor (Aurora, Northern Digital Inc.) was affixed to the distal end of the colonoscope, and pose was sampled at 40 Hz. An experienced gastroenterologist (VA) performed the simulated screening colonoscopy to capture realistic camera trajectories. In total, eight full-colon simulated screening colonoscopy videos (withdrawal phase) were recorded, each corresponding to different texture phantoms and debris. The average duration for these videos was approximately 6.5 minutes, consistent with the >6-minute withdrawal phase duration guideline issued by the American College of Gastroenterology taskforce [31].
2D-3D Registration Pipeline
The 2D-3D registration process proposed in C3VD [29] was followed. A virtual camera was traversed along the recorded real colonoscope trajectory to generate pixel-level ground truth for each video frame. This virtual camera rendered corresponding ground truth frames from the 3D model of the colon. A key challenge is determining the relative location of the phantom model in relation to the virtual camera’s trajectory. Assuming the phantom model remained stationary during the video capture, its pose can be represented by a single, rigid body transformation. To estimate this unknown transformation, we utilize a 2D-3D registration approach. This approach aligns the virtual 3D model with the 2D video frames using extracted edge maps. The registration process iteratively samples the model transform parameter space and measures alignment between edge maps from the video frames and depth maps rendered by the virtual camera.
Data Preprocessing
The recorded video frames and pose log are preprocessed before the registration step. Five keyframes were uniformly sampled from each video sequence. This number was empirically chosen to balance registration accuracy and computational efficiency in C3VD [29]. When the selected keyframes lacked informative edges (e.g., close-up of textureless regions) or contained distorted edges (e.g., lens cleaning artifacts), the closest temporally adjacent frame exhibiting clear edges was manually selected as a replacement. After selecting keyframes, preprocessing was applied to facilitate the generation of high-quality edge maps. First, specular reflections were removed using Telea inpainting [32] to reduce spurious edges. Due to the diverse range of tissue colors, textures and lighting conditions present in the phantom videos, contrast enhancement was necessary. Key frame images were first denoised with a Gaussian filter, followed by grayscale conversion. Contrast Limited Adaptive Histogram Equalization (CLAHE) [33] was then applied. Subsequently, the deep learning-based edge extraction method DexiNed[34] was utilized. The pre-trained model weights (BIPED_10.pt) available from the DexiNed repository were used for this purpose. As the DexiNed output edge images utilized an inverted format (white background, black edges), key frame edges were first inverted. Finally, non-maximal suppression and hysteresis thresholding submodules from the Canny edge detection[35] method were applied. These preprocessing steps generated binary edge maps with thin edges from the input RGB key frames.
The pose log underwent preprocessing to reduce noise. A box filter with a window size of 54 ms was applied to the data. Subsequently, the pose log was downsampled to 60 Hz. It is important to note that the pose log records the position of the robotic arm’s end-effector relative to its base, not the camera itself. Hence, a hand-eye calibration was performed to estimate the camera poses. This calibration established a transformation matrix $\mathbf{X}$ that related the relative motion of the robotic arm $\mathbf{A}$ to the relative motion of the colonoscope camera $\mathbf{B}$. This relationship was expressed by the equation $\mathbf{A}\mathbf{X}=\mathbf{X}\mathbf{B}$. The Park & Martin [36] optimization method was used to solve for the unknown transformation $\mathbf{X}$. To synchronize the video frames and the pose log, the time delay between the start of motion in the video and the pose was calculated. Synchronization was achieved by determining the relative temporal offset that maximized the correlation between the magnitude of optical flow (indicating camera motion) and the displacement recorded in the translation components of the pose. Finally, after transforming robotic arm poses into camera poses and synchronizing them with the video frames, the camera pose for each keyframe was computed by interpolation between the temporally closest camera poses.
For the simulated screening videos, temporal synchronization between video frames and EM tracker poses was established by manually identifying correspondence points, training a linear regression model (time_sensor = drift x time_video + offset) to correct clock drift, then interpolating poses to align with adjusted video frame timestamps. Clock drift arises due to the lack of hardware synchronization between the colonoscope’s video acquisition and the EM tracker, which operate on separate internal clocks. Over time, small differences in their clock frequencies accumulate, leading to misalignment between recorded timestamps. The linear model compensates for this temporal drift, enabling accurate alignment of pose data with corresponding video frames. Then, a hand-eye transform was applied to derive the camera poses from the EM tracker poses, using the same method described above.
Fisheye Camera Model
An essential feature of clinical colonoscopes is their wide FoV, which is typically around $170^{\circ}$. This feature maximizes the colon surface area visible during a colonoscopy procedure. C3VD incorporated this fisheye effect in its virtual camera model to render the pixel-wise ground truth. As in C3VD, we use the spherical omnidirectional camera model proposed in Scaramuzza et al. [37].
The spherical camera model establishes a one-to-one correspondence between a pixel $(u",v")$ on the image plane and a ray direction in 3D world space $(X,Y,Z)$. The Scaramuzza model employs a polynomial projection function to map the ideal undistorted image coordinates $(u,v)$ to the ray direction, by:
$\begin{bmatrix}X\\ Y\\ Z\end{bmatrix}=\lambda\begin{bmatrix}u\\ v\\ \alpha_{0}+\alpha_{2}\rho^{2}+\alpha_{3}\rho^{3}+\alpha_{4}\rho^{4}\end{bmatrix},$ (1)
where $\lambda$ is a scalar factor, $\alpha_{i}$ are the polynomial coefficients ($\alpha_{1}=0$), and $\rho=\sqrt{u^{2}+v^{2}}$ is the radial distance of the pixel from the image center. To account for lens-sensor misalignment and distortion, the ideal image coordinates are transformed into real distorted image pixels $(u",v")$ using the following equation:
$\begin{bmatrix}u"\\ v"\end{bmatrix}=\begin{bmatrix}c&d\\ e&1\end{bmatrix}\begin{bmatrix}u\\ v\end{bmatrix}+\begin{bmatrix}c_{x}\\ c_{y}\end{bmatrix}$ (2)
where $c,d,$ and $e$ are elements of the stretch matrix and $(c_{x},c_{y})$ correspond to center shift. These intrinsic camera parameters are computed during a camera calibration measurement. An $8\times 11$ checkerboard with a square size of 10 mm was utilized as a calibration target. The MATLAB 2023b Camera Calibration toolbox (Mathworks) was used to compute the fisheye camera intrinsics, which resulted in a mean reprojection error of 0.43 pixels. The camera intrinsics are provided in camera_intrinsics.json file.
CMAES Optimization and Edge Loss
The 2D-3D registration process aligns geometric features between the video frames and the 3D model. The optimization and loss function proposed in the original C3VD study were used. During this process, the virtual camera traversed the pre-defined camera trajectory, rendering depth maps from the 3D colon model at each keyframe pose. Within each registration iteration, the current transformation estimate (T) was applied to the 3D model. Subsequently, depth maps were rendered at camera poses corresponding to the extracted video keyframes. Edge maps were extracted from the depth images using the Canny edge detection algorithm. An edge similarity score $\mathcal{S}_{e}$ was used to quantify the similarity between the edge maps obtained from the video keyframes and the rendered depth maps. After blurring the edge maps with a Gaussian kernel, the edge similarity score between the rendered depth’s edge maps $E_{\text{Depth}}$ and the target edge maps extracted from video keyframes $E_{\text{RGB}}$ is computed as:
$\mathcal{S}_{e}=\frac{1}{\textrm{KHW}}\sum_{k=1}^{\textrm{K}}\sum_{v=1}^{\textrm{H}}\sum_{u=1}^{\textrm{W}}E_{\text{RGB}}^{k}(u,v)\cdot E_{\text{Depth}}^{k}(u,v),$ (3)
where K is the number of keyframe pairs, H and W represent the image height and width, respectively, and the dot product is applied between corresponding pixels in the edge maps. This edge similarity term was incorporated into the overall objective function used to optimize T:
$\mathbf{T}_{\mathrm{final}}=\operatorname*{argmin}_{\mathbf{T}}(1.0-\mathcal{S}_{e}),$ (4)
where $\mathbf{T}_{\mathrm{final}}$ represents the final estimated transform. The minimization was achieved using the Covariance Matrix Adaptation Evolution Strategy (CMA-ES) optimization method [38].
The initial model transform estimate $\mathbf{T}_{\mathrm{init}}$ was obtained through a manual alignment process using the 3D model-video overlay tool provided in the C3VD code repository. The CMA-ES optimization then refined this initial estimate. A population size of 100 samples was used during the optimization process. The parameter search space for the transformation was bounded within ±0.1 radians for rotations and ±7.5 millimeters for translations relative to the initial transform. Additionally, the initial step size $\sigma$ was set to 0.1. In some cases, the search space and sigma were further reduced individually for finer alignment.
The 2D-3D registration algorithm leveraged the similarity of edge features between video keyframes and rendered depth maps derived from the 3D model. Edges in the depth maps predominantly correspond to anatomical landmarks, such as haustral folds, which were partially occluded by debris in the RGB frames. To mitigate this limitation, video sequences v2 and v3 were acquired from the same phantom before and after debris, using an identical robotic camera trajectory. Given the stationary phantom orientation between acquisitions and the robotic arm’s sub-millimeter repeatability error, the model transform calculated during v2 registration was applied directly to the debris-affected v3 sequence. This approach facilitated consistent model-video alignment despite partial visual obstruction.
To synchronize the clean (v2) and debris-filled (v3) colonoscopy videos, we computed the temporal offset that maximized the correlation between their motion trajectories. Specifically, we extracted the translation components from the camera pose logs of both videos and computed their correlation while sliding one trajectory over the other. The frame offset corresponding to the highest correlation was used to align the videos. Excess frames were subsequently trimmed, resulting in paired frames of clean and debris-filled colon from the two videos.
Data Records
The dataset comprises two distinct colon shapes (c1 and c2), each segmented into seven to eight anatomical regions, with each segment further characterized by four unique textures and colors (t1, t2, t3, and t4). Overall, eight full colon phantom models were created from sixty phantom segments.
For each phantom segment, up to four video sequences were recorded. The first two videos, v1 and v2, depict a clean colon environment with variations in camera trajectories and imaging settings. The third video, v3, incorporates fecal debris, blood, foam, and other colon artifacts to simulate realistic scenarios. Notably, the camera trajectory and imaging settings for v3 are identical to that of v2. These videos are stored in the Registered Videos folder. The fourth video for each segment (v4) features phantom deformation, with stationary or linear camera motion, and always includes debris. The v4 videos can be found in the Deformation folder. A total of 184 short videos were recorded, comprising 169 rigid (v1, v2 & v3) and 15 deformation (v4) videos with a total of 74,071 frames.
During video recording, parameters such as camera trajectory, speed, and edge enhancement settings were varied to simulate diverse colonoscopy scenarios. These variations, along with simulated artifacts and challenging cases, are comprehensively documented in the C3VDv2_Data_Summary_Sheet.xlsx, explained below.
Following 2D-3D registration, ground truth maps were generated for each RGB video frame. The folder structure for each registered video is as follows:
  • RGB Frame: rgb/NNNN.png represents the raw (distorted) video frame from the Olympus CF-HQ190L video colonoscope. The black border with video metadata was cropped, resulting in an image size of 1350 x 1080 pixels. NNNN denotes the 4-digit frame number within the video.
  • Depth Frame: depth/NNNN_depth.tiff represents the depth along the camera frame’s Z-axis, clamped between 0 and 100 mm, and linearly scaled and encoded as a 16-bit grayscale image. For example, a pixel value of 16,384 corresponds to a depth of 25 mm.
  • Surface Normal Frame: normals/NNNN_normals.tiff stores the X, Y, and Z components of the surface normal vector for each surfel in the R, G, and B color channels, respectively. Components are linearly scaled from $\pm$1 to 0-65535 and encoded as a 16-bit color image. Normal vector directions are defined with respect to the camera coordinate system: +x points right, +y points down, and +z points along the viewing direction (i.e., away from the camera).
  • Optical Flow Frame: optical_flow/NNNN_flow.tiff depicts the optical flow from the current to the previous frame. X-direction motion (left to right, clamped between -20 to 20 pixels) is stored in the red channel, and Y-direction motion (up to down, clamped between -20 to 20 pixels) is stored in the green channel. Flow values are linearly scaled and encoded as a 16-bit color image.
  • Occlusion Frame: occlusions/NNNN_occlusion.tiff indicates pixels that occlude other mesh faces within 100 mm of the camera origin, assigning a value of 255 to these pixels and 0 to all others. This binary data is encoded as an 8-bit grayscale image.
  • Diffuse Frame: diffuse/NNNN_diffuse.png encodes Lambertian reflectance, computed using the dot product of the surface normal and the direction of the incident light. Reflectance values range from 0.1 to 1.0 and are linearly scaled and encoded as an 8-bit grayscale image.
  • Camera Pose: pose.txt contains each frame’s flattened homogeneous camera-to-world transformation matrix (row major order).
  • 3D Model and Coverage Map: coverage_mesh.obj stores the ground truth triangulated mesh. Texture vertices store coverage values, where vt=1 indicates an observed face, and vt=2 indicates an unobserved face.
We also include eight videos of simulated screening colonoscopy in the Simulated Screening Videos folder, comprising 95,300 frames. Each folder comprises all recorded RGB frames and a pose.txt file containing the camera poses in a frame-wise homogeneous format.
The C3VDv2 Data Summary Sheet contains the following columns:
  • Nomenclature: The video name follows the format colon_segment_texture_videonumber, corresponding to the colon shape, anatomical segment, phantom texture, and video number.
  • Relative Camera Speed: This refers to the percentage scaling factor (0-100%) of the UR3e arm’s maximum programmed velocity during motion segments. Actual trajectory velocities are non-uniform due to automatic acceleration/deceleration between waypoints.
  • Edge Enhancement: This post-processing image enhancement mode increases image sharpness according to the selected setting, ranging from 1 (smoothest) to 3 (sharpest).
  • Debris: This is a binary value of 0 for images of a clean colon and 1 for images of a colon with debris.
  • Deformation: This is a binary value with 1 indicating deformation was externally applied to the phantom during image acquisition and 0 for no deformation. This is true for v4 videos.
  • Open End Visible: Colon segment phantoms have two open ends, one of which is used for camera insertion. This binary value indicates 1 if the other end is visible in the video and 0 if the other end is not visible.
  • Tags: Contains a list of scenarios occurring in the video. The tags indicate the presence of ‘polyps’, ‘saturation’, ‘water on lens’, ‘debris on lens’, ‘water jet’, and camera motion and trajectories such as ‘fast’, ‘straight line’, ‘loop’, ‘helical’, ‘mirrored path’ (second half of the camera trajectory is the reverse of the first half), ‘zigzag’, and ‘textureless surface enface’.
  • Comments: Includes additional information such as the type of debris (e.g. blood and fecal), camera motion (e.g. enface to down-the-barrel transition), events in videos (e.g. polyp cleaning with water jet).
  • Qualitative Score: Subjective score from 1 to 3 indicating the quality of alignment between RGB frames and rendered views from the 3D model. 1 - Best Alignment, 2 - Good Alignment, 3 - Misalignment due to phantom manufacturing defects.
  • Quantitative Score: Comprises a combination of dice score (overlap) and chamfer distance (misalignment) on edges extracted from RGB frames and depth frames rendered from the 3D model. Refer to section Technical Validation for more details on this score.
Two calibration sequences are provided for geometric and photometric calibration in the Camera Calibration folder:
  • camera_calib_checker.avi: This video captures a $8\times 11$ checkerboard pattern with a square size of 10 mm.
  • camera_calib_vicalib.avi: This video features the “big_pattern” target from the Endomapper repository. The accompanying big_pattern.pdf file, when printed at 100% scale, includes a grid spacing of 5.29 mm, a large radius of 1.58 mm, and a small radius of 1.06 mm.
It is important to note that the pixel-wise depth, normals, occlusion maps, and 3D models for debris-filled colon videos (v3) are all determined from the pixel-wise ground truth for the ‘clean’ version of the colon without debris. These do not account for changes in depth, normals, or occlusion due to the addition of debris. The deformation videos (v4) and simulated screening videos do not include pixel-wise ground truth maps.
Technical Validation
Clean and Debris-Filled Colon Video Synchronization
C3VDv2 contains paired clean colon (v2) and debris-filled colon videos (v3), with identical camera trajectories and acquisition settings. To synchronize the videos, the time delay was calculated by maximizing the correlation between the translation components of the robotic arm poses. After trimming the beginnings and/or ends of the videos to account for this delay, the model transform from the clean colon video was applied to the debris-filled colon video. This reuse relies on two assumptions: the relative orientation of the model to the robotic arm remains constant between the two acquisitions, and the robotic arm has sub-millimeter pose repeatability. To validate these assumptions and the frame alignment, the similarity between paired frames was assessed using the Structural Similarity Index (SSIM). Ten random frames from the clean colon video were compared to a 31-frame window (offset +15 to -15) centered on the corresponding debris-filled colon frame. The average SSIM over the 10 frames was reported for each offset. The plot of average SSIM and standard deviation across all C3DVv2 video pairs (Figure 6a) shows a maximum at an offset of 0 frames (SSIM = 0.91), indicating maximum similarity between paired images. Importantly, the synchronization method only uses information from the camera poses, while the validation method is based only on comparing frames from the videos. This validates the synchronization method and justifies reusing the model transform for debris-filled videos.
2D-3D Registration Scoring
An edge-based 2D–3D registration method was utilized to align the 3D model with video frames by optimizing the model transformation. The registration algorithm was validated by Bobrow et al.[29] in C3VD using synthetic sequences. The proposed loss function demonstrated minimal rotational and translational error compared to alternative loss functions.
The 2D–3D registration performed well across most videos. However, we observed that the 3D model does not align optimally with the phantom in some cases, due to phantom manufacturing defects. Since the true transformation for real videos is unknown, a qualitative scoring system was implemented to assess the perceived alignment between the video and the 3D model. The qualitative alignment assessment was performed by manually reviewing video frames overlaid with their corresponding depth maps rendered from the 3D model.
Videos were rated on a scale of 1 to 3. Score 1 indicates the best alignment. Score 2 represents good alignment, where local phantom defects lead to minor misalignment. Score 3 indicates a more considerable misalignment due to manufacturing defects in the phantom.
We note that pouring silicone into a mold sometimes results in bubbles partially embedded in the casted silicone surface. After curing, extra silicone was applied to patch these bubble holes; however, this can lead to an overflow outside the holes. Consequently, when such a phantom was placed in the mold for imaging, any excess silicone will cause local deformation of the phantom shape. This was the source for most of the videos with alignment scores of 3.
Additionally, quantitative scoring was performed by evaluating the alignment of edges extracted from the 3D model and the RGB frames from the videos. For each clean colon video, the edges from all RGB and depth frames—rendered from the 3D model—were obtained using the DexiNed model. These edges were binarized, and the Dice coefficient and the Chamfer distance between the edges extracted from depth frames and the edges extracted from RGB frames were computed and averaged across all video frames. Quantitative alignment scores $Q_{\text{align}}$ were computed by combining standardized z-scores of the Dice coefficient $D_{z}$ (edge similarity) and Chamfer distance $C_{z}$ (edge displacement). The final metric is defined as:
$Q_{\text{align}}=\frac{1}{2}\left(D_{z}-C_{z}\right).$ (5)
A higher value of the quantitative alignment metric signifies better alignment. This formulation rewards high Dice scores (strong overlap) and penalizes high Chamfer distances (large misalignment). Both the Dice coefficient and the Chamfer distance are used, as the Dice coefficient is insensitive to the degree of misalignment in instances where edges do not overlap.
A comparative analysis of each video’s quantitative and qualitative metrics revealed a strong correlation, as illustrated in Figure 6b. We observed that most videos classified with a qualitative score of 1 exhibited higher quantitative alignment scores compared to those classified as 3. However, some Score 1 videos demonstrated low quantitative scores, often due to the presence of very few edges. In such cases, even minor misalignments within a subset of edges result in a notable decrease in the quantitative metric. Conversely, videos classified as Score 2 or 3, which exhibited high quantitative scores, generally maintained good alignment for most edges. However, they were categorized as Score 2 and 3 due to phantom deformations, which affected only a small subset of edges, thus allowing the quantitative score to remain high. Figure 6c shows the distribution of challenging scenarios across different scores.
Usage Notes
The dataset is available for download and preview on the project webpage. The dataset is available under the CC BY-NC-SA 4.0 license.
Discussion and Conclusion
C3VDv2 provides a comprehensive platform for the quantitative validation of 3D reconstruction algorithms such as SLAM[39, 40, 41] and its sub-components[42], and depth estimation[43, 44]. The dataset includes ground truth 3D models, camera poses, pixel-level depth, surface normals, optical flow, and occluded regions. These elements allow for precise evaluation of SLAM algorithms and their ability to identify unobserved regions during colonoscopy, with coverage maps serving as ground truth references. With the provided camera poses, C3VDv2 is also suitable for volumetric rendering based techniques such as neural radiance fields[8, 45] and Gaussian splatting[9, 46]. The wide variety of tissue texture and color in phantoms enables testing for domain generalization. Using Score 1 and 2 videos for tasks that require ground truth with high pixel-wise accuracy is recommended. In addition to camera pose estimation, Score 3 videos would be useful for training weakly supervised, semi- or self-supervised algorithms [47].
The dataset’s high-resolution 3D models and phantom-making protocols are open-sourced, enabling researchers to modify, 3D print colon molds, and cast their own phantoms for imaging. These protocols can be expanded to fabricate other tissue phantoms, such as other regions of the gastrointestinal tract like the esophagus and stomach, and to create organ structures for surgical datasets. C3VDv2 addresses scenarios where 3D reconstruction algorithms commonly fail, such as tracking loss on textureless colon surfaces, scope immersion in fluid or debris, and lens cleaning events. Additionally, the dataset includes trajectories designed to potentially debug reconstruction algorithms, including straight-line in-and-out motions, loops with identical starting and ending points, and mirrored trajectories with lens cleaning in the middle. Representative examples are shown in Figure 7. The simulated screening videos performed by an experienced gastroenterologist on complete colon phantoms offer realistic camera trajectories for validation of camera pose estimation methods. Short videos (v4) featuring simulated phantom deformations could be used for qualitative evaluation of deformable 3D reconstruction techniques.
The availability of debris-filled and corresponding clean colon videos with identical camera settings and trajectories allows researchers to assess the impact of debris on algorithm performance quantitatively. This also allows the dataset to be used for other clinically relevant tasks, such as artifact detection[48] and haustral fold detection [49] in the presence of debris. While the fecal debris has not been classified in terms of bowel prep, researchers could extend the dataset by labeling the Boston Bowel Preparatory Scale[50] to the videos to have a granular understanding of their algorithms’ performances with varying levels of debris. C3VDv2 videos and their paired ground truth frames could also be used to generate synthetic data [51, 52] for semi-supervised [53] or self-supervised learning. Since the dataset summary sheet describes the events occurring in the colonoscopy videos, the video-text pairs could aid training medical vision-language models[54]. C3VDv2 also holds potential for polyp detection and tracking. Some polyps have been strategically positioned to be occluded by haustral folds. Also, fecal debris could generate false-positive detections, presenting a challenging validation set for lesion detection algorithms.
Where pixel-level ground truth is available (v1, v2, and v3 videos), the models are constrained to be rigid and static. This contrasts with in vivo colonoscopy, where the colon is moving and deformable. Real colonic tissue has dense vasculature [55], and polyps have microscopic surface patterns, which are not replicated in the phantoms described here. These features may serve as key points for feature-tracking algorithms in real colonoscopy, but may fail in the case of phantoms.
Future datasets could include phantoms with tuned optical scattering and absorption properties to simulate realistic light-tissue interactions [56, 57, 58]. Phantoms with polyps suitable for simulated polypectomy [59] would enable the incorporation of therapeutic tools. Another future research area is generating pixel-wise ground truth and time-stamped 3D models for the deformable colon.
Code availability
The code used for the dataset is publicly available at https://github.com/DurrLab/C3VD
Acknowledgments
This work was supported in part with funding and products provided by Olympus Corporation of the Americas. Although the agreement states a Sponsored Research Agreement, Olympus is funding, but not sponsoring this research. The authors would like to thank Lauren Shepard, Jinchi Wei, Dr. Ahmed Ghazi, and Dr. Ali Uneri of the Carnegie Center for Surgical Innovation, Dr. Swaroop Vedula of Malone Center for Engineering in Healthcare, and Dr. Jianing Li, and Dr. Surya Evani of the Division of Gastroenterology & Hepatology at the Johns Hopkins Hospital for their assistance and for providing resources for data collection. The authors would also like to thank researchers at the University of North Carolina at Chapel Hill, University College London, and the University of Zaragoza for their feedback on the original C3VD dataset.
Figures & Tables
Figure 1: The C3VDv2 dataset comprises 169 colonoscopy videos with paired ground truth data, including depth, normals, occlusion, optical flow, camera pose, diffuse images, and 3D model with coverage. New enhancements in the C3VDv2 dataset videos offer improved realism, adding colonic artifacts such as fecal debris, mucous pools, blood, and foam. Challenging cases like lens cleaning, water on the lens, debris on the lens, colon deformation, and fast, non-smooth movements have been incorporated to enhance the dataset’s diversity and comprehensiveness and reduce the domain gap with in vivo colonoscopy. An additional 8 videos are provided with full colon withdrawals, and 15 videos with colon deformations.
Figure 2: Overview of the C3VDv2 data collection process. a. High-resolution 3D models of colons are digitally sculpted. b. Three-part molds are 3D printed. The high-resolution core insert (black) represents the colon lumen, while the outer shells (white) define the external contours of the phantom. c. The phantom casting process involves (1) painting silicone layer-by-layer onto the insert mold with the vasculature hand-painted on intermediate layers. The painted insert is then placed in the outer shells, and (2) silicone is poured through mold inlets to fill the space between the insert and outer shell, d. resulting in the final colon phantom. e. The imaging protocol involves rigidly attaching the colonoscope tip to a robotic arm for precise, repeatable, and pose-tracked maneuvers. Phantoms are secured inside their outer shells to maintain shape. f. 2D-3D registration is performed to align the 3D model with the 2D RGB video frames. The loss function is optimized based on the overlap of the corresponding edge maps. g. After alignment, a virtual camera is navigated along the camera trajectory to render pixel-wise ground truth frames.
Figure 3: Cross-section of Colon models: High-resolution 3D models were digitally sculpted referencing anatomical images. For positional reference, magnified views of lesions and Ileocecal valves are shown. The colors represent the distinct colon segments used for making phantoms. Quantitative details such as polyp sizes are provided in Table 2 and 3.
Figure 4: Sample frames from ‘clean’ colon sequences (Video v1 & Video v2): Four phantoms with varied tissue colors and textures were cast for each colon model. Paired ground truth maps are available for all RGB frames.
Figure 5: Sample frames from debris-filled colon sequences (Video v3): C3VDv2 bridges the gap between phantom datasets and real-world colonoscopy by including colonic artifacts like fecal debris, mucous pools, blood, foam, en-face textureless wall views, and water on lens, which often pose difficulties for 3D reconstruction algorithms. The frame-wise ground truth depth, normal, occlusion maps, and 3D model align with the ‘clean’ colon ground truth and do not take into account the additional debris.
Figure 6: Technical validation results: a) To verify the synchronization between the clean (v2) and debris-filled (v3) videos, the structural similarity index (SSIM) was computed across a window of frames with offsets. Maximum SSIM at offset 0 indicates that clean & debris-filled colon videos are well synchronized. b) The alignment of the RGB video frames and the depth maps was qualitatively and quantitatively evaluated. The quantitative score comprises a combination of dice score and chamfer distance, while the qualitative score was given by visual inspection of RGB video frames overlaid on rendered depth maps (1 - best, 3 - least). The qualitative scores agree with the quantitative scores. c) Histogram of challenging scenarios with respect to qualitative score. This indicates the presence of realistic scenarios, even in well-aligned videos with scores of 1 & 2.
Figure 7: Video snapshots of challenging scenarios: The C3VDv2 dataset features a wide range of challenging scenarios such as fast non-smooth motion, scope dipping to debris, lens cleaning, water on the lens, water jet, exploratory helical motion, loops, and en face to down the barrel transitions. Evaluating 3D reconstruction algorithms on C3VDv2 will provide a more representative assessment of their performance in real-world clinical settings. Green boxes indicate the same image.
Table 1: Comparison of existing monocular colonoscopy datasets
Camera type Res. FOV Tissue Type 3D GT Pose 3D GT Type Debris Blood Foam Water on Lens Deformation Frames Dataset
USB SD Narrow Phantom None None 23,935 Fulton et al [60]
Rendered SD Narrow Virtual CT Depth 16,016 Rau et al [20]
Rendered SD Narrow Virtual CT Depth 23,421 SimCol3D [21, 22]
Rendered SD/HD Narrow Virtual CT None 14,181 Dongdong et al [28]
USB SD/HD Wide Ex-vivo Porcine OS 3D Model 39,406 EndoSLAM[24]
Pill Cam SD Wide Ex-vivo Porcine None None 3,294 EndoSLAM [24]
Colonoscope HD Wide Phantom CT 3D Model 12,250 EndoSLAM [24]
Rendered SD Narrow Virtual CT Depth 21,887 EndoSLAM [24]
Colonoscope HD Wide In vivo Human None None 24 hrs video EndoMapper [25]
Colonoscope HD Wide Phantom Sculpted Depth, Normals, Flow, 3D Model 10,015 C3VD[29]
Colonoscope HD Wide Phantom Sculpted Depth, Normals, Flow, 3D Model, Diffuse 67,886 C3VDv2 (Proposed)
Colonoscope HD Wide Phantom Sculpted Undeformed 3D Model 6,185 C3VDv2 (Proposed)
Table 2: Ground truth 3D colon model attributes for Colon 1
Segment Location Length (mm) Object number Type Paris Classification Major Axis (mm)
A Cecum 125 1 Adenoma Ip 2.8 & 1.4
2 Hyperplastic Is 4.3
3 Ileocecal Valve - 24.8
B Ascending 164 - - - -
C Transverse 1 132 4 Adenoma IIa + IIs 10
D Transverse 2 160 - - - -
E Descending 149 5 Adenoma Is 3
F Sigmoid 1 150 6 Adenoma IIa 6.7
G Sigmoid 2 108 - - - -
H Rectum 167 - - - -
Table 3: Ground truth 3D colon model attributes for Colon 2
Segment Location Length (mm) Object number Type Paris Classification Major Axis (mm)
A Cecum 83 1 Ileocecal valve - 15.3
B Ascending 160 2 3 Adenoma Adenoma Ip IIa 5.7 11.4
C Transverse 1 148 4 Hyperplastic Is 5.1
D Transverse 2 100 5 Hyperplastic IIa 4.4
E Descending 170 6 Adenoma Is 3.9
F Sigmoid 94 7 8 Adenoma Hyperplastic Is IIa 3.3 11.5
G Rectum 105 - - - -